A chemist needs to prepare 32 ounces of a 14% hydrochloric acid solution. Find the amount of 16% solution and the amount of 8% solution he should mix to get this solution.
I am so confused!

Answer by TimothyLamb(4379)

---
x + y = 32
8x + 16y = 14*32
---
put the system of linear equations into standard form
---
x + y = 32
8x + 16y = 448
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 8 oz of 8%
y= 24 oz of 16%
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php